OpenStudy (anonymous):

Amanda and Devoney both worked hard over the summer. Together they earned a total of $450. Amanda earned $25 less than Devoney. Write a system of equations for the situation. Use a for the amount Amanda earned and d for the amount Devoney earned.

OpenStudy (anonymous):

@eric_d @RANE @thadyoung @UsukiDoll @paki @arabpride @arnessa_xoxo @shrutipande9 @shamim @dakari22 @Firejay5 @ganeshie8 @gnrfan @hockeychick23 @jagatuba @Compassionate @bibby @nikvist @mustafa2014

OpenStudy (anonymous):

this is why i hate math because you will never need to do this in real life to find out how much someone made..... :/ also, sorry i suck at this type of math sorry

OpenStudy (anonymous):

its ok

OpenStudy (rane):

Total is = $450 Devoney= \[x \] Amanda's earning = \[x-25\] try writing an equation now.

OpenStudy (arabpride):

Well start by using what you already know :3

OpenStudy (hockeychick23):

ok so lets make the equation have only 1 variable, we'll use d to symbolize devoney. D=d A=d-25 so lets combine the two equations: d+d-25=450 2d-25=450 Now lets add 25 to both sides 2d= 475 Divide both sides by 2. 237.50= d Devoney made 237.50 Amanda earned 25 less so 237.50-25=212.50 Amanda earned 212.50 and Devoney earned 237.50

OpenStudy (arabpride):

The total being $450 you know you will have something that looks like $450 = __ __ __

OpenStudy (arabpride):

I'd say your equation would look something like: 450 = (d - 25) + d where (d - 25) = a :3

OpenStudy (anonymous):

@hockeychick23 Thats very helpful thanks :)

OpenStudy (hockeychick23):

no problem! :)

OpenStudy (arabpride):

3,000 questions answered. Lel cx

OpenStudy (rane):

lol @arabpride ~ CONGRATS !!!! ~

OpenStudy (arabpride):

Thanks!!! :D @RANE